# Sample-Efficient Differentially Private Fine-Tuning via Gradient Matrix Denoising

### v1

@@ 5 Limitations and Future Work
One major limitation of our method is that it is does not always produces the best performance as well as the fastest convergence. In some of the experiments, the baseline method achieves slightly better final accuracy than our method. This is specially puzzling because of the dominantly positive improvement in cosine similarity between the denoised and noisy gradients with respect to the clipped gradients. This calls for further investigation to understand why this happens, and how to fix it.
@@ 6 Reproducibility Statement
Appendix A Finite dimensional derivation of random matrix theory results
In the usual random matrix theory literature, the results in Shabalin & Nobel (2013); Donoho et al. (2018); Gavish & Donoho (2014) are stated in the asymptotic regime, where the matrix dimensions grow to infinity and the noise variance may scale with the dimensions. In this section we want to state those results in their original form, and explain the derivation of equations 3, 4, and 5 from their asymptotic forms.
The setup in Shabalin & Nobel (2013); Donoho et al. (2018) is as follows. We have a sequence of matrices $\bm{X}_{n}\in\mathbb{R}^{m_{n}\times n}$ with $m_{n}/n\to\beta$ as $n\to\infty$. The rank of the signal matrix is fixed, i.e. $\mathrm{rank}(\bm{X}_{n})=r$ for all $n$. The singular values of the signal matrix are fixed, i.e. the non-zero singular values of $\bm{X}_{n}$ are $\lambda_{1}>\lambda_{2}>\ldots>\lambda_{r}>0$ for all $n$. Then, we add a noise matrix with i.i.d. entries from $\mathcal{N}(0,1/n)$ to get the noisy matrix. In these settings, the results in Shabalin & Nobel (2013); Gavish & Donoho (2014) state that
@@ Appendix A Finite dimensional derivation of random matrix theory results
where $y_{n,i}$ is the $i$-th singular value of the noisy matrix. If we want to change this into the finite dimensional form, we can start form a noise matrix with i.i.d. entries from $\mathcal{N}(0,\sigma^{2})$ instead of $\mathcal{N}(0,1/n)$. Then, if we work with the matrix $\frac{Y}{\sigma\sqrt{n}}$, then, the new noise matrix will have the desired distribution. Using the equation 9 for the matrix $\frac{Y}{\sigma\sqrt{n}}$, and substituting $\beta=m/n$, we get to the equation 3. Similar arguments can be used to derive equations 4 and 5 from their asymptotic forms in Shabalin & Nobel (2013).
Appendix B Why norm correction is needed?
We want to show in this section that why improving cosine similarity of one component of a noisy vector to the signal may not necassiryly improve the overall cosine similarity, and why norm correction helps prevent this issue.
B.1 Setup and Definitions
Let vectors be partitioned into two components:
$a=(a_{1},a_{2}),\quad b=(b_{1},b_{2}),$
with $a_{1},b_{1}\in\mathbb{R}^{d_{1}}$ and $a_{2},b_{2}\in\mathbb{R}^{d_{2}}$.
We define the following concepts for vectors partitioned into two components as above.
Overall cosine similarity: The cosine similarity between $a$ and $b$ is
$\cos(a,b)=\frac{a_{1}\cdot b_{1}+a_{2}\cdot b_{2}}{\|a\|\;\|b\|}.$
Block cosine similarity: The cosine similarity between the second blocks is
$\cos(a_{2},b_{2})=\frac{a_{2}\cdot b_{2}}{\|a_{2}\|\;\|b_{2}\|}.$
Improvement of block similarity: Given a modified block $a_{2}^{\prime}$, we say that $a_{2}^{\prime}$ improves the similarity of block $2$ with respect to $b_{2}$ if
$\cos(a_{2}^{\prime},b_{2})>\cos(a_{2},b_{2}).$
We define the modified full vector as $a^{\prime}=(a_{1},a_{2}^{\prime})$.
B.2 Improving a block is not sufficient globally
It is possible that $\cos(a_{2}^{\prime},b_{2})>\cos(a_{2},b_{2})$ but
$\cos(a^{\prime},b)<\cos(a,b).$
@@ B.2 Improving a block is not sufficient globally
We can write
$\cos(a,b)=\frac{a_{1}\cdot b_{1}+\|a_{2}\|\|b_{2}\|\cos\theta}{\sqrt{\|a_{1}\|^{2}+\|a_{2}\|^{2}}\;\|b\|},$
$\cos(a^{\prime},b)=\frac{a_{1}\cdot b_{1}+\|a_{2}^{\prime}\|\|b_{2}\|\cos\theta^{\prime}}{\sqrt{\|a_{1}\|^{2}+\|a_{2}^{\prime}\|^{2}}\;\|b\|},$
where $\theta,\theta^{\prime}$ are the angles between $a_{2},b_{2}$ and $a_{2}^{\prime},b_{2}$.
Even if $\cos\theta^{\prime}>\cos\theta$, the numerator can decrease when $\|a_{2}^{\prime}\|<\|a_{2}\|$. Since the denominator also changes with $\|a_{2}^{\prime}\|$, it is possible that $\cos(a^{\prime},b)<\cos(a,b)$. Explicit counterexamples confirm this. ∎
B.3 Equal block norm guarantees improvement
If $\|a_{2}^{\prime}\|=\|a_{2}\|$ and $\cos(a_{2}^{\prime},b_{2})>\cos(a_{2},b_{2})$, then
$\cos(a^{\prime},b)>\cos(a,b).$
Proof.
Let $r=\|a_{2}\|=\|a_{2}^{\prime}\|$ and $s=\|b_{2}\|$. Then
$a_{2}\cdot b_{2}=rs\cos\theta,\qquad a_{2}^{\prime}\cdot b_{2}=rs\cos\theta^{\prime},\qquad\cos\theta^{\prime}>\cos\theta.$
Numerators:
$N=a_{1}\cdot b_{1}+rs\cos\theta,\qquad N^{\prime}=a_{1}\cdot b_{1}+rs\cos\theta^{\prime},$
so $N^{\prime}>N$.
Denominators:
$\|a\|=\sqrt{\|a_{1}\|^{2}+r^{2}}=\|a^{\prime}\|,\qquad\|b\|\ \text{fixed}.$
Therefore
$\cos(a^{\prime},b)=\frac{N^{\prime}}{\|a\|\|b\|}>\frac{N}{\|a\|\|b\|}=\cos(a,b).$
Appendix C Use of LLMs

### v2

@@ 5 Limitations and Future Work
One major limitation of our method is that it does not always produce the best final performance decpite the fastest convergence. In some experiments, the baseline achieves slightly higher final accuracy than our method. This is particularly puzzling given the consistently positive improvement in cosine similarity between the denoised and noisy gradients relative to the clipped gradients. Further investigation is needed to understand this discrepancy and to identify possible remedies.
@@ 6 Reproducibility Statement
Appendix A Random Matrix Theory Background
We have included some additional figures and derivations for the random matrix theory results used in this paper.
A.1 Low rank structure in gradients of large language models
A.2 Finite dimensional derivation of random matrix theory results
In the usual random matrix theory literature, the results in Shabalin and Nobel (2013); Donoho et al. (2018); Gavish and Donoho (2014) are stated in the asymptotic regime, where the matrix dimensions grow to infinity and the noise variance may scale with the dimensions. In this section we want to state those results in their original form, and explain the derivation of equations 3, 4, and 5 from their asymptotic forms.
The setup in Shabalin and Nobel (2013); Donoho et al. (2018) is as follows. We have a sequence of matrices $\bm{X}_{n}\in\mathbb{R}^{m_{n}\times n}$ with $m_{n}/n\to\beta$ as $n\to\infty$. The rank of the signal matrix is fixed, i.e. $\mathrm{rank}(\bm{X}_{n})=r$ for all $n$. The singular values of the signal matrix are fixed, i.e. the non-zero singular values of $\bm{X}_{n}$ are $\lambda_{1}>\lambda_{2}>\ldots>\lambda_{r}>0$ for all $n$. Then, we add a noise matrix with i.i.d. entries from $\mathcal{N}(0,1/n)$ to get the noisy matrix. In these settings, the results in Shabalin and Nobel (2013); Gavish and Donoho (2014) state that
@@ A.2 Finite dimensional derivation of random matrix theory results
where $y_{n,i}$ is the $i$-th singular value of the noisy matrix. If we want to change this into the finite dimensional form, we can start form a noise matrix with i.i.d. entries from $\mathcal{N}(0,\sigma^{2})$ instead of $\mathcal{N}(0,1/n)$. Then, if we work with the matrix $\frac{Y}{\sigma\sqrt{n}}$, then, the new noise matrix will have the desired distribution. Using the equation 10 for the matrix $\frac{Y}{\sigma\sqrt{n}}$, and substituting $\beta=m/n$, we get to the equation 3. Similar arguments can be used to derive equations 4 and 5 from their asymptotic forms in Shabalin and Nobel (2013).
Appendix B Why denoising is needed?
As discussed in introductory paragraph of section 3, the slowdown in convergence of DP-SGD can be attributed to two main factors: the per-sample gradient clipping and the addition of noise. In this appendix, we present more empirical evidence to support this claim in figure 6.
Figure 6: Comparison between validation loss curves of DP-SGD with and without noise addition. The experiments are conducted on the SST-2 and MNLI dataset using the RoBERTa family models. The results clearly indicate that the addition of noise significantly slows down the convergence of the training process compared to the scenario where only gradient clipping is applied. We also see significant gap between the two curves in terms of final loss achieved, for some of the model/dataset pairs .
Appendix C Additional results on why threshold is needed.
Here we include additional hyperparameter sweeps for $\kappa$ on different model/dataset pairs. The results are presented in figures 7, 8, and 9. Also, scatter plot of layer improvement vs $\frac{\lambda_{1}}{\sigma(\sqrt{n}+\sqrt{m})}$ for different layer dimensionality is presented in figure 10 to further justify our choice of threshold.
Figure 7: Effect of different values of $\kappa$ on the per-layer improvements when training RoBERTa large model on SST-2 dataset. The Setting are similar to what is described in 4.
Figure 8: Effect of different values of $\kappa$ on the per-layer improvements when training RoBERTa base model on MNLI dataset. The Setting are similar to what is described in 4.
Figure 9: Effect of different values of $\kappa$ on the per-layer improvements when training RoBERTa large model on MNLI dataset. The Setting are similar to what is described in 4.
Appendix D Additional results on Why norm correction is needed.
In this appendix, we give a proof of theorem 1, as well as additional experimental results showing the abalation of norm correction in our denoising function.
D.1 Proof of Theorem 1
@@
Since the blocks are disjoint and $\|y_{i}\|=\|y^{\prime}_{i}\|$ for all $i$, we can write
$\langle y,x\rangle\;=\;\sum_{i=1}^{c}\langle y_{i},x_{i}\rangle\qquad\text{and}\qquad\|y\|^{2}\;=\;\sum_{i=1}^{c}\|y_{i}\|^{2}\;=\;\sum_{i=1}^{c}\|y^{\prime}_{i}\|^{2}\;=\;\|y^{\prime}\|^{2}.$
For each block, cosine similarity is
$\cos(y_{i},x_{i})=\frac{\langle y_{i},x_{i}\rangle}{\|y_{i}\|\|x_{i}\|}.$
Assumption (i) and the norm equality (ii) imply
$\langle y_{i},x_{i}\rangle=\|y_{i}\|\|x_{i}\|\,\cos(y_{i},x_{i})\;\leq\;\|y^{\prime}_{i}\|\|x_{i}\|\,\cos(y^{\prime}_{i},x_{i})=\langle y^{\prime}_{i},x_{i}\rangle.$
Summing over all $i$ gives $\langle y,x\rangle\leq\langle y^{\prime},x\rangle$. Using the equality of global norms, we obtain
$\cos(y,x)=\frac{\langle y,x\rangle}{\|y\|\,\|x\|}\;\leq\;\frac{\langle y^{\prime},x\rangle}{\|y^{\prime}\|\,\|x\|}=\cos(y^{\prime},x).$
Thus $\cos(y,x)\leq\cos(y^{\prime},x)$, as claimed. ∎
D.2 Additional norm correction abalation results
For the sake of completeness, we calarify what we mean by the abalation. Instead of the equation 9, we use the denoised gradient without the norm correction, i.e.,
$\mathrm{Denoise_{\textrm{unscaled}}}(\tilde{\bm{g}}\!\mid_{\bm{W}})=\begin{cases}\tilde{\bm{g}}\!\mid_{\bm{W}},&\text{if }\lambda_{1}(\tilde{\bm{g}}\!\mid_{\bm{W}})<\kappa\,\sigma(\sqrt{n}+\sqrt{m})\\[6.45831pt] \hat{\bm{g}}_{\mathrm{optimal}},&\text{otherwise}\end{cases}$
Other than the 3, we also have the results for training both RoBERTa base and large models on MNLI dataset, as well as RoBERTa large model on SST-2 dataset. These results are presented in figures 6. It is evident from these results that without the norm correction, the improvement is not consistent, and can even be negative in some cases. Also, it is evident that with the norm correction, the improvement is consistently positive. This shows the importance of the norm correction step in our denoising function.
Figure 12: Comparison of whole gradient improvement when using norm correction and not using it. The Settings are similar to what is described in 4. It is evident that using norm correction results in more consistent improvement.
Appendix E Additional language generation experiments
In this appendix, we present additional experiments on language generation tasks, specifically on the E2E benchmark (Novikova et al., 2017) and DART (Nan et al., 2021) datasets. We used the models from Qwen3 (Team, 2025) and Llama3.2 (Grattafiori et al., 2024) family. For training these models, we used LoRA (Hu et al., 2022) with rank $32$. We used a learning rate of 2e42e-4, batch size of $64$, and clipping norm of $1.0$. The privacy budget $\epsilon$ was set to $5.4$ for all experiments, and the noise multiplier was calculated using the privacy accountant from Gopi et al. (2021) to ensure a total privacy loss of $5.4$ at $400$ training steps. For generating text during evaluation, we used the same setting as in Yu et al. (2021). For the denoising, we used the same $\kappa=1.02$ as in all other experiments, showing the robustness of our method to this hyperparameter. Similar to experiments on the GLUE benchmark datasets, we see significant improvements in performance after a limited number of training steps. Here, we present the results after $50$ training steps to highlight the sample efficiency of our denoising method. In table 4, we can see that out of the 50 different model/dataset/metric combinations, the denoised models outperformed the baseline in 49 cases, demonstrating the effectiveness of our denoising method in enhancing model performance on language generation tasks under differential privacy constraints.
The final performance of the models after $400$ steps are reported in table 5. In the metrics after 400 steps, out of the 50 different model/dataset/metric combinations, the denoised models outperformed the baseline in 35 cases, showing that our method not only improves sample efficiency but also leads to better final performance in many scenarios. This, however, is still an area for further investigation to understand the cases where the baseline outperforms the denoised models at the later steps.
Dataset Model Size Variant BLEU ROUGE-L METEOR NIST CIDEr
E2E Qwen 0.6B baseline 23.17 46.78 0.509 2.34 0.67
denoised 37.35 53.52 0.628 4.60 1.11
E2E Qwen 1.7B baseline 36.01 53.57 0.609 4.64 1.08
denoised 36.95 54.42 0.624 4.71 1.14
E2E Qwen 4B baseline 40.00 55.19 0.659 4.83 1.27
denoised 40.88 55.74 0.663 5.03 1.36
E2E Llama 1B baseline 22.15 44.24 0.476 2.59 0.47
denoised 36.33 52.86 0.626 4.62 1.13
E2E Llama 3B baseline 10.96 25.22 0.268 0.93 0.22
denoised 26.33 44.36 0.516 3.48 0.64
DART Qwen 0.6B baseline 14.66 33.44 0.323 0.91 0.59
denoised 23.58 46.46 0.462 2.87 0.84
DART Qwen 1.7B baseline 21.77 46.98 0.483 4.23 0.86
denoised 33.29 51.95 0.579 5.26 1.26
DART Qwen 4B baseline 21.74 44.47 0.522 3.84 0.86
denoised 21.78 46.98 0.484 4.22 0.86
DART Llama 1B baseline 8.88 37.38 0.368 2.86 0.41
denoised 13.81 43.12 0.454 3.68 0.73
DART Llama 3B baseline 6.12 30.08 0.319 2.13 0.26
denoised 9.47 37.33 0.375 2.95 0.47
Table 4: Comparison of baseline vs. denoised models on E2E and DART datasets after 50 steps. Bold indicates the better value within each pair.
Dataset Model Size Variant BLEU ROUGE-L METEOR NIST CIDEr
E2E Qwen 0.6B baseline 38.77 53.57 0.661 4.83 1.17
denoised 38.94 55.07 0.668 4.87 1.20
E2E Qwen 1.7B baseline 40.00 55.70 0.661 4.91 1.22
denoised 40.05 55.73 0.670 4.92 1.23
E2E Qwen 4B baseline 40.58 56.47 0.678 5.00 1.32
denoised 40.44 56.33 0.676 4.99 1.28
E2E Llama 1B baseline 37.06 54.94 0.629 4.65 1.21
denoised 39.70 55.29 0.662 4.93 1.30
E2E Llama 3B baseline 39.02 54.68 0.659 4.89 1.25
denoised 39.64 54.83 0.682 4.91 1.29
DART Qwen 0.6B baseline 25.99 50.44 0.595 3.38 1.01
denoised 25.43 50.23 0.510 3.37 1.02
DART Qwen 1.7B baseline 31.07 52.63 0.572 4.77 1.26
denoised 30.98 52.89 0.581 4.93 1.26
DART Qwen 4B baseline 29.70 52.22 0.571 4.63 1.22
denoised 34.82 54.71 0.594 5.29 1.38
DART Llama 1B baseline 20.03 46.40 0.466 3.91 0.79
denoised 19.41 46.85 0.479 3.95 0.81
DART Llama 3B baseline 23.56 48.01 0.512 4.60 1.00
denoised 21.51 44.25 0.461 3.80 0.87
Table 5: Comparison of baseline vs. denoised models on E2E and DART datasets. Bold indicates the better value within each pair.
Appendix F Use of LLMs